# Worse than Random? An Embarrassingly Simple Probing Evaluation of Large Multimodal Models in Medical VQA

### v1

@@ Abstract
Large Multimodal Models (LMMs) have shown remarkable progress in the field of medical Visual Question Answering (Med-VQA), achieving high accuracy on existing benchmarks. However, their reliability under robust evaluation is questionable. This study reveals that state-of-the-art models, when subjected to simple probing evaluation, perform worse than random guessing on medical diagnosis questions. To address this critical evaluation problem, we introduce the Probing Evaluation for Medical Diagnosis (ProbMed) dataset to rigorously assess LMM performance in medical imaging through probing evaluation and procedural diagnosis. Particularly, probing evaluation features pairing original questions with negation questions with hallucinated attributes, while procedural diagnosis requires reasoning across various diagnostic dimensions for each image, including modality recognition, organ identification, clinical findings, abnormalities, and positional grounding. Our evaluation reveals that top-performing models like GPT-4V and Gemini Pro perform worse than random guessing on specialized diagnostic questions, indicating significant limitations in handling fine-grained medical inquiries. Besides, models like LLaVA-Med struggle even with more general questions, and results from CheXagent demonstrate the transferability of expertise across different modalities of the same organ, showing that specialized domain knowledge is still crucial for improving performance. This study underscores the urgent need for more robust evaluation to ensure the reliability of LMMs in critical fields like medical diagnosis, and current LMMs are still far from applicable to those fields. https://github.com/eric-ai-lab/ProbMed
Figure 1: Accuracy of four LMMs on two types of specialized questions in medical diagnoses, with and without adversarial pairs. The significant drop in accuracy with adversarial pairs highlights the models’ unreliability in handling medical diagnoses.
@@ 1 Introduction
Foundation models, such as large language models (LLMs) (OpenAI, 2023a; Touvron et al., 2023; Jiang et al., 2023; Anil et al., 2023; Chung et al., 2022) and large multimodal models (LMMs) (OpenAI, 2023b; Team et al., 2023; Li et al., 2023b; Liu et al., 2023a; Chen et al., 2023a), have demonstrated impressive capabilities in understanding complex visual and text inputs, generating human-like language, and achieving high accuracy on various benchmarks. The integration of these foundation models into real-life medical practice holds immense potential given their advanced computational capabilities (Wu et al., 2023; Yang et al., 2023) and promising progress on existing medical Visual Question Answering (Med-VQA) benchmarks (Lau et al., 2018; Liu et al., 2021; He et al., 2020; Zhang et al., 2023). As we stand on the precipice of integrating these models into critical decision-making domains, one natural question appears: how much can we trust these models in real-world scenarios, such as medicine and healthcare, where the stakes are high?
Before discussing the reliability of LMMs in critical domains like Med-VQA, we must first address a fundamental question: Are we evaluating LMMs correctly? To address this question, we introduce a simple yet effective probing evaluation method that exposes the weaknesses of LMMs by creating simple binary questions with hallucination pairs over existing benchmarks. An example is shown in Figure 2. Despite the high accuracy reported on current Med-VQA tasks, our study reveals a significant vulnerability in LMMs when faced with adversarial questioning, as illustrated in Figure 1. The performance drops observed are alarming: even advanced models like GPT-4V and Gemini Pro perform worse than random guessing, with an average decrease of 42% across the tested models.
@@
Our systematic evaluation of seven state-of-the-art LMMs on ProbMed revealed several critical insights. First, even the best-performing models, such as GPT-4V and Gemini Pro, performed close to random guessing on specialized diagnostic categories like Condition/Finding and Position, highlighting their limitations in handling fine-grained medical inquiries. Second, introducing adversarial pairs significantly reduced the accuracy of all models, with LLaVA-Med’s performance dropping by up to 57% and Gemini Pro’s accuracy decreasing by 25% in ProbMed. These findings emphasize the importance of adversarial testing in Med-VQA to uncover model weaknesses. Third, the CheXagent model, which was exclusively trained on chest X-rays, demonstrated that specialized domain knowledge is crucial. It showed that expertise gained on one particular organ could be transferable to another modality of the same organ in a zero-shot manner, highlighting the value of domain-specific training for improving model performance.
In summary, our work highlights significant gaps in the reliability of LMMs for medical diagnosis despite their impressive performance on current existing general domain benchmarks. The insights from ProbMed underscore the urgent need for robust evaluation methodologies to ensure the accuracy and reliability of LMMs in real-world medical applications. This research inspires the development of more trustworthy AI systems in healthcare and beyond, ultimately contributing to better diagnostic outcomes and patient care.
@@ 2 Related Work
2.1 Large Multimodal Models in the Medical Domain
The advancements in Large Multimodal Models (LMMs) have significantly enhanced the understanding and generation of medical content that integrates both visual and linguistic elements. Notable models include GPT-4V (OpenAI, 2023b), Gemini Pro (Team et al., 2023), LLaVA (Liu et al., 2023a, b), and MiniGPT-v2 (Chen et al., 2023a). The scalability and exceptional performance of these large foundation models have driven their application in the biomedical field.
Further progress has been made in fine-tuning general-domain LMMs for the biomedical field, resulting in specialized models like LLaVA-Med (Li et al., 2023a), MedBLIP (Chen et al., 2023b), and Med-Flamingo (Moor et al., 2023). Despite the promising results from these domain-specific LMMs, there is ongoing exploration into training smaller multimodal models to address specific clinical needs. For instance, models like LLaVA-RAD (Chaves et al., 2024) and CheXagent (Chen et al., 2024) have been developed for chest X-ray interpretation, aiming to bridge competency gaps in radiology tasks. As we move towards integrating these models into critical decision-making processes, it becomes imperative to assess their reliability in high-stakes environments like healthcare and medicine.
2.2 Medical Visual Question Answering
@@
Automated curation methods have been developed to address scalability. PathVQA (He et al., 2020) uses CoreNLP [^1] https://stanfordnlp.github.io/CoreNLP tools, and PMC-VQA (Zhang et al., 2023) employs generative models to create larger datasets. However, these methods often sacrifice fine-grained question categories, and some require additionally trained models for question filtering. ProbMed, as shown in Table 1, stands out by providing large-scale benchmarks and enabling categorical accuracy assessments across various diagnostic dimensions for each image, including modality recognition, organ recognition, clinical findings identification, and positional grounding. ProbMed uniquely incorporates adversarial negation pairs for each question-answer pair to ensure diagnostic specificity and reliability, setting it apart from existing benchmarks.
Different evaluation methods are employed for assessing LMMs, including closed-ended VQA, multiple choice VQA, and open-ended generation tasks such as captioning and report generation. Open-ended VQA and report generation are typically considered more challenging and harder to evaluate, often requiring human or model evaluation alongside automated lexical similarity metrics like ROUGE-L and BLEU-4. Recent works (Wang et al., 2024; Zheng et al., 2024) argue that multiple-choice questions may not be ideal due to inherent "selection bias." In our work, we choose a relatively easy-to-evaluate method: closed-ended VQA augmented with adversarial evaluation methods featuring hallucinated attributes. By requiring the model to accurately distinguish relevant features, we enhance the reliability of the evaluation process. This method allows for clear and definitive assessments, improving the overall robustness of our findings in medical contexts.
Table 1: Comparison ProbMed with a test set of existing medical VQA datasets, demonstrating our dataset’s difference from existing benchmarks. For SLAKE, only the English subset is considered for head-to-head comparison with existing benchmarks.
Dataset Images Questions
Question
Category
Procedural
Diagnosis
Adversarial
Pairs
VQA-RAD (Lau et al., 2018) 0.2k 0.4k
SLAKE (Liu et al., 2021) 0.09k 1k
PathVQA (He et al., 2020) 0.8k 6.7k
PMC-VQA (Zhang et al., 2023) 50k 400k
ProbMed (Ours) 6.3k 57k
3 ProbMed: Probing Evaluation for Medical Diagnosis
In this section, we design two evaluation principles and present a comprehensive analysis of state-of-the-art LMMs for Med-VQA using the created ProbMed dataset to address two research questions:
@@ 3 ProbMed: Probing Evaluation for Medical Diagnosis
Our primary goal is to rigorously evaluate these models’ readiness for real-life diagnostic tasks, particularly under adversarial conditions. Despite their high accuracy on existing benchmarks, the models struggle with simple probing evaluation. ProbMed is designed to expose these vulnerabilities and provide a more reliable assessment of model performance in real-world scenarios.
@@ 3.1 Probing Evaluation with Adversarial Pairs
One of the main motivations behind ProbMed is to assess the models’ ability to accurately distinguish between relevant and irrelevant features. ProbMed pairs original questions with negation questions containing hallucinated attributes. This method challenges the model robustness by requiring them to identify true conditions while disregarding false, hallucinated ones. For instance, a question about a specific finding is paired with a negated question featuring a different, non-existent finding to test if the model can exclusively identify the true finding.
@@ 3.2 Procedural Diagnosis
To ensure a comprehensive evaluation, ProbMed includes questions that require reasoning across multiple diagnostic dimensions for each image. These dimensions include modality recognition, organ identification, clinical findings, abnormalities, and positional reasoning. This multifaceted approach assesses a model’s diagnostic capabilities beyond simple question-answer pairs, requiring it to integrate various pieces of information to form a coherent diagnostic picture.
3.3 Data Curation
@@
As illustrated in Figure 3, ProbMed draws from two comprehensive biomedical datasets MedICaT and ChestX-ray14 to compile a diverse set of 6,303 images. MedICaT (Subramanian et al., 2020) contains 217k image-caption pairs from 131k open access biomedical papers. From this dataset, we selected 4,543 image-caption pairs focusing on a single organ and modality, with clear indications of normal or abnormal conditions. These images span three modalities (X-ray, MRI, and CT scan) and four organs (abdomen, brain, chest, and spine). ChestX-ray14 (Wang et al., 2017) comprises 112k frontal-view X-ray images from 30k unique patients, including 880 images with abnormalities marked by bounding boxes. We selected 1,760 images, balanced between healthy and abnormal cases, with disease labels and bounding box annotations.
To create high-quality, balanced question-answer pairs, we generated metadata for each image. For MedICaT images, GPT-4 (OpenAI, 2023a) was used to analyze captions, identify abnormalities, and extract positional descriptions using few-shot prompting. For ChestX-ray14 images, a positional reasoning module generated textual descriptors of abnormalities based on bounding boxes and image sizes. The standardized metadata across both subsets included unique condition names and positional descriptions for each organ-modality combination as candidates for later constructing adversarial question-answers.
Using the metadata, we generated a diverse set of high-quality questions for each image, covering various diagnostic dimensions. Ground-truth questions were created based on available metadata, with "yes" answers. For each ground-truth question, we also created a corresponding adversarial question by selecting random adversarial entities and assigning "no" answers. This process included selecting alternative organs, modalities, and hallucinated conditions/findings. As shown in Table 2, this process resulted in a total of 57,132 question-answer pairs, averaging 9 pairs per image.
@@
4 Experimental Analysis
We conducted a systematic evaluation and comprehensive analysis using the ProbMed dataset on seven state-of-the-art LMMs to identify their strengths and weaknesses in real-life imaging diagnostics. Apart from proprietary GPT-4V (OpenAI, 2023b) and Gemini Pro (Team et al., 2023), we selected five open-source models (7B parameters version), spanning across general models including LLaVA-v1 (Liu et al., 2023b), LLaVA-v1.6 (Liu et al., 2023a), MiniGPT-v2 (Chen et al., 2023a) and specialized models including LLaVA-Med (Li et al., 2023a) and CheXagent (Chen et al., 2024). These models were chosen based on their computational cost, efficiency, and inference speed, making them practical for integration into medical practice. For categorical accuracy, we only considered a hit if the model correctly answered all questions within the category for an image (see Table 19). This means the model must correctly identify all existing entities and disregard all hallucinated entities.
4.1 Is Current Evaluation of LMMs for Med-VQA Reliable?
To address this first research question, we introduced adversarial pairs to the evaluation process to test the model’s robustness and reliability. This strategy ensures that models must validate the absence of certain characteristics or findings rather than simply acknowledge existing conditions, thereby enhancing diagnostic specificity and reliability. To demonstrate the necessity of adversarial pairs for achieving valid and trustworthy accuracy scores in Med-VQA, we conduct an experimental analysis on the test set of an existing medical dataset, VQA-RAD (Lau et al., 2018), in addition to ProbMed.
4.1.1 Probing Evaluation with Adversarial Pairs in VQA-RAD
To construct challenging adversarial questions for an given image, ideally, we need full control over the ground truth information and a set of confusing candidates, as provided in ProbMed. However, since VQA-RAD (Lau et al., 2018) provides finalized question-answer pairs without metadata, we could only construct adversarial pairs for 118 test instances where the answer is "yes" out of 272 closed-ended question-answers pairs within its test set. Each adversarial pair was manually created such that, based on the limited information from the original question-answer pair, the answer to the adversarial question had to be negated. This process resulted in 236 question-answer pairs in total. The adversarial questions in this subset are less challenging than those in ProbMed, as they often involve a simple semantic negation of the original question due to limited information.
These results, as shown in Table 3, reveal the significant impact of adversarial pairs on model performance. Although the original accuracy appears very high for some underperforming models, the accuracy drops drastically after balancing the subset with adversarial pairs: 6.78% for GPT-4V and 16.95% for Gemini Pro, with an average decrease of 42.7% across the tested models.
Table 3: Model accuracy on the VQA-RAD test subset before and after introducing adversarial pairs. The table demonstrates the significant drop in accuracy across various models when adversarial pairs are added, highlighting the models’ vulnerabilities to adversarial questions. The percentage decrease in accuracy is noted in parentheses.
Models
Accuracy (%) without
Adversarial Pairs

### v5

@@ Abstract
Large Multimodal Models (LMMs) have demonstrated impressive performance on existing medical Visual Question Answering (Med-VQA) benchmarks. However, high reported accuracy does not necessarily reflect their true diagnostic reliability in clinical settings. This study reveals that state-of-the-art models perform worse than random guessing on medical diagnosis questions when subjected to simple Probing Evaluation for Medical Diagnosis (ProbMed). ProbMed challenges models through probing evaluation and procedural diagnosis. Particularly, probing evaluation features pairing ground-truth questions with adversarial counterparts that feature negated and hallucinated attributes, while procedural diagnosis requires reasoning across various dimensions for each image, including modality recognition, organ identification, clinical findings, abnormalities, and positional grounding. Our evaluation reveals that even top-performing models like GPT-4o, GPT-4V, and Gemini Pro perform worse than random guessing on specialized diagnostic questions, indicating significant limitations in handling fine-grained medical inquiries. Furthermore, our ablation study on open-source models (e.g., LLaVA, LLaVA-Med, and Med-Flamingo) identifies poor visual understanding as a primary bottleneck—a limitation that can be partially mitigated by incorporating visual descriptions generated by GPT-4o, resulting in an average performance improvement of 9.44%. These findings underscore the urgent need for more robust evaluation methods and domain-specific expertise to ensure the reliability of LMMs in high-stakes medical applications.
Worse than Random? An Embarrassingly Simple Probing Evaluation of Large Multimodal Models in Medical VQA
Qianqi Yan1, Xuehai He1, Xiang Yue2, Xin Eric Wang1 1University of California, Santa Cruz, 2Carnegie Mellon University
Figure 1: Accuracy of six LMMs on two types of specialized questions in medical diagnoses, with and without adversarial pairs. The significant drop in accuracy with adversarial pairs highlights the models’ unreliability in handling medical diagnoses with our probing method.
@@ 1 Introduction
Foundation models, such as large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Jiang et al., 2023; Anil et al., 2023; Chung et al., 2024) and large multimodal models (LMMs) (OpenAI, 2024, 2023; Team et al., 2024; Li et al., 2023; Liu et al., 2023a; Chen et al., 2023), have demonstrated impressive capabilities in understanding complex visual and text inputs, generating human-like language, and achieving high accuracy on various benchmarks. The integration of these foundation models into real-life medical practice holds immense potential given their advanced computational capabilities (Wu et al., 2023a; Yang et al., 2023) and promising progress on existing medical Visual Question Answering (Med-VQA) benchmarks (Lau et al., 2018; Liu et al., 2021; He et al., 2020; Zhang et al., 2023). As we stand on the precipice of integrating these models into critical decision-making domains, one natural question appears: how much can we trust these models in real-world scenarios, such as medicine and healthcare, where the stakes are high?
Before discussing the reliability of LMMs in critical domains like Med-VQA, we must first address a fundamental question: Are we evaluating LMMs correctly? To address this question, we introduce a simple yet effective probing evaluation method that exposes the weaknesses of LMMs by creating binary questions with hallucination pairs over existing benchmarks. An example is shown in Figure 2. Despite the high accuracy reported on current Med-VQA tasks, our study reveals a significant vulnerability in LMMs when faced with adversarial questioning, as illustrated in Figure 1. The observed performance drops are alarming: even advanced models like GPT-4o, GPT-4V, and Gemini Pro perform worse than random guessing, with an average decrease of 27.78% across the tested models.
@@
Our systematic evaluation of twelve state-of-the-art LMMs on ProbMed revealed several critical insights. First, even the best-performing models, such as GPT-4o and Gemini Pro, performed close to random guessing on specialized diagnostic categories like Condition/Finding and Position, highlighting their limitations in handling fine-grained medical inquiries. Second, introducing adversarial pairs significantly reduced the accuracy of all models, with LLaVA-Med-v1.5’s performance dropping by up to 28.22% and GPT-4o’s accuracy decreasing by 20.71% in ProbMed. These findings emphasize the importance of adversarial testing in Med-VQA to uncover model weaknesses. Third, by incorporating chain-of-thought reasoning and adding visual descriptions generated by GPT-4o, we observe substantial improvements in model performance, suggesting that poor visual understanding is a critical bottleneck. The results indicate that augmenting these models with more accurate visual information could significantly improve their ability to handle complex medical tasks. Moreover, the CheXagent model, which was exclusively trained on chest X-rays, demonstrated that specialized domain knowledge is crucial. It showed that expertise gained on one particular organ could be transferable to another modality of the same organ in a zero-shot manner, highlighting the value of domain-specific training for improving model performance.
In summary, our work highlights significant gaps in the reliability of LMMs for medical diagnosis despite their impressive performance on existing general domain benchmarks. The insights from ProbMed underscore the urgent need for robust evaluation methodologies to ensure the accuracy and reliability of LMMs in real-world medical applications. Our findings also suggest that poor visual understanding is a key limitation for open-source models, which can be mitigated by incorporating chain-of-thought reasoning and accurate visual descriptions. This research inspires the development of more trustworthy AI systems in healthcare and beyond, ultimately contributing to better diagnostic outcomes and patient care.
@@ 2 Related Work
Large Multimodal Models in the Medical Domain
The advancements in Large Multimodal Models (LMMs) have significantly enhanced the understanding and generation of medical content that integrates both visual and linguistic elements. Notable models include GPT-4o (OpenAI, 2024), GPT-4V (OpenAI, 2023), Gemini Pro (Team et al., 2024), LLaVA (Liu et al., 2023a, 2024), and MiniGPT-v2 (Chen et al., 2023). The scalability and exceptional performance of these large foundation models have driven their application in the biomedical field.
Further progress has been made in fine-tuning general-domain LMMs for the biomedical field, resulting in specialized models like BiomedGPT (Zhang et al., 2024), LLaVA-Med (Li et al., 2024), Med-Flamingo (Moor et al., 2023), MedBLIP (Chen and Hong, 2024), RadFM (Wu et al., 2023b) and MedVInT (Zhang et al., 2023). Despite the promising results from these domain-specific LMMs, ongoing exploration exists into training smaller multimodal models to address specific clinical needs. For instance, models like LLaVA-RAD (Chaves et al., 2024) and CheXagent (Chen et al., 2024) have been developed for chest X-ray interpretation, aiming to bridge competency gaps in radiology tasks.
Comprehensive surveys of LLMs for healthcare highlight the progress, applications, and challenges in deploying LLMs in clinical settings (He et al., 2023; Zhou et al., 2024; Peng et al., 2023). Task-specific evaluations  (Yan et al., 2023; Liu et al., 2023b) underline the potential and challenges of LMMs in the medical domain. As we move towards integrating these models into critical decision-making processes, it becomes imperative to assess their reliability in high-stakes environments like healthcare and medicine.
Medical Visual Question Answering
@@
Automated curation methods have been developed to address scalability. PathVQA (He et al., 2020) uses CoreNLP [^1] https://stanfordnlp.github.io/CoreNLP tools, and PMC-VQA (Zhang et al., 2023) employs generative models to create larger datasets. However, these methods often sacrifice fine-grained question categories, and some require additionally trained models for question filtering.
Different evaluation methods are employed for assessing LMMs, including closed-ended VQA, multiple choice VQA, and open-ended generation tasks such as captioning and report generation. Open-ended VQA and report generation are typically considered more challenging and harder to evaluate, often requiring human or model evaluation alongside automated lexical similarity metrics like ROUGE-L and BLEU-4. Recent works (Wang et al., 2024; Zheng et al., 2024; Zong et al., 2024) argue that multiple-choice questions may not be ideal due to inherent selection bias and permutation sensitivity. In our work, we choose a relatively easy-to-evaluate method: closed-ended VQA augmented with adversarial evaluation methods featuring hallucinated attributes. By requiring the model to accurately distinguish relevant features, we enhance the reliability of the evaluation process. This method allows for clear and definitive assessments, improving the overall robustness of our findings in medical contexts.
3 Probing Evaluation for Medical Diagnosis
In this section, we introduce two complementary evaluation strategies that rigorously assess state-of-the-art LMMs for Med-VQA. Our approach is designed to answer the following research questions:
@@ 3 Probing Evaluation for Medical Diagnosis
Despite current high accuracy, we find that the models struggle with simple probing evaluation on existing benchmarks. Our evaluation framework, ProbMed (Probing Evaluation for Medical Diagnosis), is built on adversarial testing and multifaceted diagnostic reasoning to further expose these vulnerabilities and provide a thorough analysis of model performance. We also explore enhancements through chain-of-thought reasoning and the incorporation of external visual descriptions (e.g., from GPT-4o) to address the noted limitations of open-sourced models.
@@ 3.1 Probing Evaluation with Adversarial Pairs
A core element of our framework is the use of adversarial pairs to test model robustness. For each image, we generate pairs of: ground-truth questions that query the presence of a specific entity (e.g., a particular clinical finding) with corresponding adversarial questions that introduce a negated or hallucinated attribute (e.g., a non-existent finding or an alternative organ). This pairing challenges the model to discern between actual diagnostic features and spurious details, thereby revealing its ability—or inability—to filter out irrelevant or misleading information. The performance drop observed under adversarial conditions highlights the fragility of current evaluation protocols and motivates the need for more robust assessment methods.
@@ 3.2 Procedural Diagnosis
Beyond binary question-answering, ProbMed incorporates procedural diagnosis to evaluate the models’ diagnostic reasoning. Each image is associated with questions spanning multiple diagnostic dimensions, including Modality Recognition: Identifying the imaging technique (e.g., X-ray, MRI, CT). Organ Identification: Determining the anatomical region under investigation. Clinical Findings and Abnormalities: Detecting abnormal conditions. Positional Reasoning: Localizing findings spatially within the image. This multifaceted evaluation framework requires models to integrate diverse pieces of information for each test sample, thereby providing a more comprehensive measure of their diagnostic capabilities.
3.3 Data Filtering and Curation
ProbMed is curated from two widely recognized biomedical datasets: MedICaT and ChestX-ray14. The data curation process, summarized in Figure 3, involves the following steps:
Image Selection:
@@ 3.3 Data Filtering and Curation
From MedICaT (Subramanian et al., 2020), we extracted 4,543 image-caption pairs that focus on a single organ and modality with clear indications of normal or abnormal conditions. From ChestX-ray14 (Wang et al., 2017), we selected 1,760 frontal-view X-ray images balanced between healthy and abnormal cases, including those with bounding box annotations for abnormalities.
Metadata Generation:
For each image, we generate a unified metadata record:
$D_{i}=\{\mathtt{mod}_{i},\mathtt{organ}_{i},\{\mathtt{condition}_{j},\mathtt{pos}_{j}\}_{j}^{n_{i}}\}$
where $\mathtt{mod}_{i}$ and $\mathtt{organ}_{i}$ denote the imaging modality and anatomical region, respectively, and each $\{\mathtt{condition}_{j},\mathtt{pos}_{j}\}$ pair represents a detected clinical finding and its positional description. For MedICaT, GPT-4 is leveraged via few-shot prompting to extract abnormality details and positional cues from captions. For ChestX-ray14, a dedicated positional reasoning module generates descriptive text based on bounding box coordinates.
3.4 Evaluation Protocol
For each diagnostic entity in the metadata, we automatically generate a ground-truth question $Q_{i}$, asking the model to confirm the presence of that specific entity. An adversarial question $Q_{i}^{\prime}$, constructed by randomly selecting an alternative or hallucinated entity (e.g., an incorrect organ or false condition) and expecting a “no” response.
Crucially, our accuracy metric is defined in a strict manner: an entity is considered correctly predicted only if the model provides the correct answer for both $Q_{i}$ and $Q_{i}^{\prime}$. In other words, if a model answers “yes” to both questions for a given entity, it is deemed incorrect rather than receiving partial credit. For images containing more than one $\{\mathtt{condition}_{j},\mathtt{pos}_{j}\}$ pair, the accuracy under the Condition/Finding and Position category is computed as the average accuracy over all $n_{i}$ pairs—there is no bonus for partial correctness. This evaluation setup ensures that only unambiguous, fully correct responses are counted as hits, highlighting the models’ true diagnostic reliability. (See Appendix B for detailed statistics on the number of questions in each category.)
3.5 Expert Study
To validate the reliability of our metadata and the corresponding question-answer pairs, we conducted an expert verification study. Two medical experts independently reviewed 100 randomly sampled metadata entries out of 6,303 from ProbMed, as well as the 1,090 QA pairs corresponding to those metadata entries. The review process yielded an average accuracy of 94.00% for the metadata and 97.79% for the QA pairs. This rigorous validation underscores the quality and thorough curation of the ProbMed dataset. In addition, we conducted a study using the expert-verified data on GPT-4o to verify the robustness of our probing evaluation framework to dataset errors, detailed in Appendix C.
As of statistics reported in Table 1, our data curation process produced a total of 57,132 question-answer pairs (averaging 9 pairs per image), spanning a comprehensive set of diagnostic dimensions. These high-quality, balanced pairs provide a robust foundation for evaluating model performance.
@@ 3.5 Expert Study
4 Experiments and Analysis
Table 2: Model accuracy on the VQA-RAD test subset and ProbMed with adversarial pairs. Accuracy is reported in two ways: (1) averaged across individual questions in a pair and (2) requiring both the ground truth and adversarial questions for the same image to be answered correctly. The drop in accuracy across models demonstrates their vulnerability to adversarial questions, with percentage decreases shown in parentheses.
Models VQA-RAD ProbMed
Averaged
Accuracy (%)